# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 8. Flag manifolds
Flag manifolds find applications in data science in recent years, see [30, 27, 24, 17] for surveys. Additional background for flag manifolds could be found in [30, 21, 19]. The simplified parallel transport condition in section 6 is satisfied if $\beta=-1$, or $\alpha=\frac{1}{2}$, but not in general. We provide a summary relevant to the computation of geodesics and parallel transport.
@@
With $n>d$, for a matrix $Y\in\mathrm{St}_{n,d}$, consider a partition of columns of $Y$ to $p$ blocks of sizes $(d_{1},\cdots,d_{p})$, $Y=\begin{bmatrix}Y_{1}|\cdots|Y_{p}\end{bmatrix}$, $Y_{i}\in\mathbb{R}^{n\times d_{i}}$. Consider the block-diagonal group $\mathrm{K}_{\mathfrak{a}}=\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p}))\subset\mathrm{SO}(d)$, whose element $U=\operatorname{diag}(U_{1},\cdots U_{p})$ acts on $Y$ by right multiplication, $YU=\begin{bmatrix}Y_{1}U_{1}|\cdots|Y_{p}U_{p}\end{bmatrix}$ ($\mathrm{S}$ means $\det(U)=1$).
@@
A flag manifold is the quotient of $\mathrm{St}_{n,d}$ by this action [30, section 4.2]. It is also typically expressed as a quotient of $\mathrm{SO}(n)$, allowing us to leverage our framework. Set $d_{p+1}=n-d$, the flag manifold $\mathrm{St}_{n,d}/\mathrm{K}_{\mathfrak{a}}$ will be denoted by $\operatorname{Flag}(\vec{d})$, where $\vec{d}=(d_{1},\cdots,d_{p},d_{p+1})$. We can also consider $\operatorname{Flag}(\vec{d})$ as a quotient of $\mathrm{SO}(n)$ by the diagonal blockdiagonal subgroup $\mathrm{K}=\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p})\times\operatorname{O}(d_{p+1}))$. Going forward, we will use the term flag diagonal blocks to this block structure. Thus, in our setting, $\mathrm{G}=\mathrm{SO}(n),\mathfrak{g}=\mathfrak{so}(n),\mathfrak{a}=\mathfrak{so}(d)$ identified with the top left $d\times d$ block, $\mathfrak{k}=\mathfrak{so}(d_{1})\times\cdots\times\mathfrak{so}(d_{p})\times\mathfrak{so}(d_{p+1})$. Hence, $\mathfrak{k}$ is a block diagonal Lie algebra of antisymmetric matrices, while its complement $\mathfrak{m}$ is the space of antisymmetric matrices with zero flag diagonal blocks. $\mathfrak{k}_{\mathfrak{a}}$ is the Lie algebra of $\mathrm{K}_{\mathfrak{a}}$ above, while $\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$ is the bottom $\mathfrak{so}(n-d)$. We show a diagram of the spaces in table 1 for $p=2$.
$n$ldelim{3* $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$
$\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$
$\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}=\mathfrak{a}_{{}_{\top}}$
$\scriptstyle d_{1}$ $\scriptstyle d_{2}$ $\scriptstyle d_{3}=n-d$
$\to$ $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}}$ $\mathfrak{k}_{\mathfrak{a}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\mathfrak{m}_{\mathfrak{a}_{{}_{\Join}}}$ $\scriptstyle d=d_{1}+d_{2}$
Table 1. Arrangement of subspaces and subalgebras for a flag manifold $\operatorname{Flag}(d_{1},d_{2},d_{3})$ as a quotient of $\mathrm{SO}(n)$ (left) or $\mathrm{St}_{n,d}$ (right) for $p=2$. The tangent space at $\operatorname{I}_{n}$, represented by an $n\times n$ matrix in $\mathfrak{so}(n)$ is divided to cells corresponding to subspaces and subalgebras. The top-left $2\times 2$ cells are in $\mathfrak{a}$, the bottom-right cell is $\mathfrak{a}_{{}_{\top}}$. The quotient is by the block-diagonal group $\mathrm{K}$ corresponding to $\mathfrak{k}=\mathfrak{k}_{\mathfrak{a}}\oplus\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$. The metric is constructed using $\mathfrak{a}$. The image is the horizontal space at $\operatorname{I}_{n,d}\in\mathrm{St}_{n,d}$, with the last column block removed and the $\mathfrak{k}_{\mathfrak{a}}$ cells set to zero.
Under the metric with parameter $\vec{\beta}=(-\frac{1}{2},\alpha)$ in section 5, we note if $\mathtt{a}_{\mathfrak{a}}\in\mathfrak{a}$ with zero flag diagonal blocks (thus $\mathtt{a}_{\mathfrak{a}}\in\mathfrak{a}\cap\mathfrak{m}$), and $U(t)=\exp(t(2\alpha-1)\mathtt{a}_{\mathfrak{a}})$, then $U(t)^{\mathsf{T}}\mathfrak{k}U(t)$ may not be in $\mathfrak{k}$, and the simplified condition for parallel transport is not expected to apply in general. However, it applies for $\alpha=\frac{1}{2}$ where $U(t)=\operatorname{I}_{n}$. This is the bi-invariant metric on $\mathrm{SO}(n)$, the formula for $\mathrm{P}_{\mathtt{a}}$ in eq. 39 reduces to $\mathrm{P}_{\mathtt{a}}w=\frac{1}{2}[w,\mathtt{a}]_{\mathfrak{m}}$.
@@
To compute in the more efficient Stiefel coordinates, we proceed as before. In Stiefel coordinates, the Levi-Civita connection is given by [21, eq. 6.4].
@@
(53) $\nabla^{\mathcal{H}}_{\mathtt{Z}}\mathtt{W}=\mathrm{D}_{\mathtt{Z}}\mathtt{W}+Y(\mathtt{Z}^{\mathsf{T}}\mathtt{W})_{\mathrm{symf}}+(1-\alpha)(\operatorname{I}_{n}-YY^{\mathsf{T}})(\mathtt{Z}\mathtt{W}^{\mathsf{T}}+\mathtt{W}\mathtt{Z}^{\mathsf{T}})Y.$
@@
Here, for a matrix $A\in\mathbb{R}^{d\times d}$, $A_{\mathrm{symf}}=\frac{1}{2}(A+A^{\mathsf{T}})+\frac{1}{2}(A-A^{\mathsf{T}})_{\mathfrak{k}_{\mathfrak{a}}}$, the projection to $\mathfrak{k}_{\mathfrak{a}}$ is the operation extracting the diagonal blocks corresponding to $\mathfrak{so}(d_{i})$, $i=1\cdots p$. Thus, we symmetrize $A$ but add back the antisymmetric parts of the flag diagonal blocks, hence $A_{\mathrm{symf}}$ means symmetrize $A$ but leaving the flag diagonal blocks unchanged. This accounts for the difference between the full tangent space of a Stiefel manifold and the horizontal space of a flag manifold, as seen from eq. 38. This is proved in [21]. Alternatively, we can project the formula eq. 38, similar to the Stiefel case in [20].
@@
Thus, given an initial horizontal vector $\xi$ in $T_{Y}\mathrm{St}_{n,p}$, we can form the vector subspace $\mathcal{S}\subset\mathbb{R}^{n\times d}$ with columns spaned by columns of $Y$ and $\xi$. Decompose $\xi=YA+QR$ with $Q\in\mathbb{R}^{n\times k}$, then $\mathcal{M}_{\mathcal{S}}$ is isometric via the map $V\mapsto F_{YQ}V$ to $\mathrm{St}_{d+k,k}$, where $F_{YQ}=\begin{bmatrix}Y|Q\end{bmatrix}$ as before. The group $\mathrm{K}_{\mathfrak{a}}=\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p}))$ operates on $\mathcal{M}_{\mathcal{S}}$ and on $\mathrm{St}_{d+k,k}$. By associativity $F_{YQ}(VU)=(F_{YQ}V)U$ if $U\in\mathrm{SO}(d)\supset\mathrm{K}_{\mathfrak{a}}$, thus horizontal spaces map to horizontal spaces. Therefore, flag geodesics and transports between these two spaces correspond, and we obtain geodesic formulas in reduced form. Similarly, we can map parallel transports with initial horizontal vectors in $\mathcal{M}_{\mathcal{S}}$, where we can use the simplified transport formula when $\alpha=\frac{1}{2}$.
@@
More specifically, let $\eta$ be a horizontal vector to be transported, we decompose $\eta=\eta_{\mathcal{S}}+\eta_{\mathcal{S}^{o}}$, where $\eta_{\mathcal{S}^{o}}$ is horizontal and normal to both $Y$ and $\mathcal{S}$. Similar to the Stiefel case, the transport of $\eta_{\mathcal{S}^{o}}$ is given by eq. 49. This is proved exactly like in the Stiefel case, using eq. 53, with $\eta_{\mathcal{S}^{o}}=(\operatorname{I}_{n}-YY^{\mathsf{T}})\eta-QQ^{\mathsf{T}}\eta$, then $\gamma^{\mathsf{T}}\Delta_{\mathcal{S}^{o}}=\dot{\gamma}^{\mathsf{T}}\Delta_{\mathcal{S}^{o}}=0$, and the rest follows by solving a linear ODE explicitly. Thus,
@@
For $\alpha=\frac{1}{2}$, at $Y\in\mathrm{St}_{n,p}$, with $\xi=YA+QR$ is a horizontal vector as in section 7, the parallel transport of a horizontal vector $\eta\in T_{Y}\mathrm{St}_{n,p}$ of $\operatorname{Flag}(\vec{d})$ lifted to $\mathrm{St}_{n,p}$ by
@@
With $\mathtt{a}=\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}$ and $w=\begin{bmatrix}w_{a}&-w_{r}^{\mathsf{T}}\\ w_{r}&0\end{bmatrix}$ then $\mathrm{P}$ is given by
(55) $\mathrm{P}(w)=\mathrm{P}_{\mathtt{a}}(w):=\begin{bmatrix}((w_{a}A+R^{\mathsf{T}}w_{r})_{\mathrm{skew}})_{\mathfrak{m}}\\ \frac{1}{2}(w_{r}A-Rw_{a})\end{bmatrix}.$
For the proof, we simplify eq. 52. The subscript $\mathfrak{m}$ on the top $d\times d$ block means setting the flag diagonal blocks to zero.
@@ 9. Numerical studies
We implemented the formulas in Python. The tables and graphs are produced from Google colab workbooks in the folder examples in [22] on default free session setting with two Xenon CPUs and $13$GB of RAM.

### v2

@@ 9. Flag manifolds with the canonical metric
Flag manifolds have found applications in data science in recent years, see [44, 40, 34, 23] for surveys. Additional background for flag manifolds could be found in [44, 31, 29]. As we will see, the naturally reductive space construction in section 7 works for the original bi-invariant (canonical) metric. We provide here a summary relevant to the computation.
@@
With $n>d$, for a matrix $Y\in\mathrm{St}_{n,d}$, consider a partition of columns of $Y$ to $p$ blocks of sizes $(d_{1},\cdots,d_{p})$, $Y=\begin{bmatrix}Y_{1}|\cdots|Y_{p}\end{bmatrix}$, $Y_{i}\in\mathbb{R}^{n\times d_{i}}$. Consider the block diagonal group $\mathcal{K}_{St}=\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p})\subset\operatorname{O}(d)$, whose element $U=\operatorname{diag}(U_{1},\cdots U_{p})$ acts on $Y$ by right multiplication, $YU=\begin{bmatrix}Y_{1}U_{1}|\cdots|Y_{p}U_{p}\end{bmatrix}$.
@@
A flag manifold is the quotient of $\mathrm{St}_{n,d}$ by this action [44, section 4.2]. It is also typically expressed as a quotient of $\mathrm{SO}(n)$, allowing us to leverage our framework. Set $d_{p+1}=n-d$, the flag manifold $\mathrm{St}_{n,d}/\mathcal{K}_{St}$ will be denoted by $\operatorname{Flag}(\vec{d})$, where $\vec{d}=(d_{1},\cdots,d_{p},d_{p+1})$. We can also consider $\operatorname{Flag}(\vec{d})$ as a quotient of $\mathrm{SO}(n)$ by the block diagonal subgroup $\mathrm{S}(\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p})\times\operatorname{O}(d_{p+1}))$ [44] ($\mathrm{S}$ denotes the subgroup of determinant $1$).
Going forward, we will use the term flag diagonal blocks for this block structure. Set $\mathcal{G}=\mathrm{SO}(n),\mathfrak{g}=\mathfrak{so}(n)$ and $\mathfrak{k}=\mathfrak{so}(d_{1})\times\cdots\times\mathfrak{so}(d_{p})\times\mathfrak{so}(d_{p+1})$. Hence, $\mathfrak{k}$ is a block diagonal Lie algebra of antisymmetric matrices, while its complement $\mathfrak{m}$ is the space of antisymmetric matrices with zero flag diagonal blocks. Any nontrivial block subalgebra $\mathfrak{a}$ would overlap $\mathfrak{k}$, thus, to use the result of section 7, we will take $\mathfrak{a}=\{0\}$ and consider the naturally reductive manifold $\tilde{\mathcal{G}}/\tilde{\mathcal{K}}$ with $\tilde{\mathcal{G}}=\mathcal{G}$ and $\tilde{\mathcal{K}}=\mathcal{K}$. Note that this metric corresponds to $\alpha=\frac{1}{2}$ in sections 6 and 8, with $\beta_{0}=-\frac{1}{2}$ and $\mathcal{B}=\frac{1}{2}\operatorname{I}_{\mathfrak{a}}$.
@@
To compute in the more efficient Stiefel coordinates, we proceed as before. In Stiefel coordinates, the Levi-Civita connection is given by [31, eq. 6.4].
@@
(63) $\nabla^{\mathcal{H}}_{\mathtt{Z}}\mathtt{W}=\mathrm{D}_{\mathtt{Z}}\mathtt{W}+Y(\mathtt{Z}^{\mathsf{T}}\mathtt{W})_{\mathrm{symf}}+\frac{1}{2}(\operatorname{I}_{n}-YY^{\mathsf{T}})(\mathtt{Z}\mathtt{W}^{\mathsf{T}}+\mathtt{W}\mathtt{Z}^{\mathsf{T}})Y.$
@@
Here, we use $\alpha=\frac{1}{2}$ and for a matrix $A\in\mathbb{R}^{d\times d}$, $A_{\mathrm{symf}}=\frac{1}{2}(A+A^{\mathsf{T}})+\frac{1}{2}\mathcal{P}_{\mathfrak{k}}(A-A^{\mathsf{T}})$, the projection to $\mathfrak{k}$ is the operation extracting the diagonal blocks corresponding to $\mathfrak{so}(d_{i})$, $i=1\cdots p$. Thus, we symmetrize $A$ but add back the antisymmetric parts of the flag diagonal blocks, hence $A_{\mathrm{symf}}$ means symmetrizing $A$ but leaving the flag diagonal blocks unchanged. This accounts for the difference between the full tangent space of a Stiefel manifold and the horizontal space of a flag manifold (see the extra $\mathcal{P}_{\mathfrak{k}}$ term in eq. 49). This is proved in [31], but we can also derive it from eq. 49.
@@
Given an initial horizontal vector $\xi$ in $T_{Y}\mathrm{St}_{n,p}$, we can form the vector subspace $\mathcal{S}\subset\mathbb{R}^{n\times d}$ with columns spanned by columns of $Y$ and $\xi$. Decompose $\xi=YA+QR$ with $Q\in\mathbb{R}^{n\times k}$, then $\mathcal{M}_{\mathcal{S}}$ is isometric via the map $V\mapsto F_{YQ}V$ to $\mathrm{St}_{d+k,k}$, where $F_{YQ}=\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}$ as before. The group $\mathcal{K}_{St}=\operatorname{O}(d_{1})\times\cdots\times\operatorname{O}(d_{p})$ operates on $\mathcal{M}_{\mathcal{S}}$ and on $\mathrm{St}_{d+k,k}$. By associativity $F_{YQ}(VU)=(F_{YQ}V)U$ if $U\in\operatorname{O}(d)\supset\mathcal{K}_{St}$, thus horizontal spaces map to horizontal spaces. Therefore, flag geodesics and transports between these two spaces correspond, and we obtain geodesic formulas in Stiefel coordinates. Similarly, we can map parallel transports with initial horizontal vectors in $\mathcal{M}_{\mathcal{S}}$.
@@
More specifically, let $\eta$ be a horizontal vector to be transported, we decompose $\eta=\eta_{\mathcal{S}}+\eta_{\mathcal{S}^{o}}$, where $\eta_{\mathcal{S}^{o}}$ horizontal and normal to both $Y$ and $\mathcal{S}$. Similar to the Stiefel case, the transport of $\eta_{\mathcal{S}^{o}}$ is given by lemma 5. This is proved exactly like in the Stiefel case, using eq. 63, with $\eta_{\mathcal{S}^{o}}=(\operatorname{I}_{n}-YY^{\mathsf{T}})\eta-QQ^{\mathsf{T}}\eta$, then $\gamma^{\mathsf{T}}\Delta_{\mathcal{S}^{o}}=\dot{\gamma}^{\mathsf{T}}\Delta_{\mathcal{S}^{o}}=0$, and the rest follows by solving a linear ODE explicitly. Thus,
@@
With the metric eq. 54 for $\alpha=\frac{1}{2}$, at $Y\in\mathrm{St}_{n,p}$, let $\xi=YA+QR$ be a horizontal vector as in section 8, the parallel transport of a horizontal vector $\eta\in T_{Y}\mathrm{St}_{n,p}$ of $\operatorname{Flag}(\vec{d})$ lifts to $\mathrm{St}_{n,p}$ by
@@
With $\mathtt{V}=\begin{bmatrix}A&-R^{\mathsf{T}}\\ R&0\end{bmatrix}$ and $\mathtt{W}=\begin{bmatrix}\mathtt{W}_{a}&-\mathtt{W}_{r}^{\mathsf{T}}\\ \mathtt{W}_{r}&0\end{bmatrix}$ then $\mathrm{P}$ is given by
(65) $\mathrm{P}(\mathtt{W})=\mathrm{P}_{\mathtt{V}}(\mathtt{W}):=\begin{bmatrix}\mathcal{P}_{\mathfrak{m}}(\mathtt{W}_{a}A+R^{\mathsf{T}}\mathtt{W}_{r})_{\mathrm{skew}}\\ \frac{1}{2}(\mathtt{W}_{r}A-R\mathtt{W}_{a})\end{bmatrix}$.
The proof is by substituting in eqs. 50 and 51 then take the first $d$ columns as in the proof of eq. 62. The operator $\mathcal{P}_{\mathfrak{m}}$ in eq. 65 sets the flag diagonal blocks to zero.
Remark 4.
We can equip the quotient $\mathrm{St}_{n,d}/\mathcal{K}_{St}$ with the deformation metric on $\mathrm{St}_{n,d}$ for $\alpha\neq\frac{1}{2}$, giving the flag manifold a different metric. It can be shown by substitution that in this case, while the Levi-Civita connection and the geodesics are of similar formats, the part of the parallel transport equation corresponding to the exponential action becomes the solution of an ODE with nonconstant coefficients.
@@ 10. Numerical studies
We implemented the formulas in Python. The tables and graphs are produced from Google colab workbooks in the folder examples in [32] on free sessions with two Xenon CPUs and 1313GB of RAM.
In algorithm 1, we describe the parallel transport algorithm for Stiefel and flag manifolds (with $\alpha=\frac{1}{2}$ for flag manifolds). In the decomposition $\xi=YA+QR$, $A=Y^{\mathsf{T}}\xi$, while $Q$ and $R$ could be computed by the SVD or QR decomposition of $\xi-YY^{\mathsf{T}}\xi$. We call algorithm 2 to calculate $W$ in the parallel transport. As indicated in appendix A, instead of computing $\operatorname{expv}(\mathrm{P}_{AR,\mathcal{M}},B)$ directly (the operator denotes the map $\mathrm{P}_{AR}$ in either eq. 62 or eq. 65 based on $\mathcal{M}$), we define the function $\mathsf{S}_{c}:\begin{bmatrix}T\\ B\end{bmatrix}\mapsto\begin{bmatrix}cT\\ B\end{bmatrix}$ scaling the top block of a matrix, then compute the exponential of $\mathrm{P}^{bal}_{AR,\mathcal{M}}=\mathsf{S}_{\alpha^{\frac{1}{2}}}\circ\mathrm{P}_{AR,\mathcal{M}}\circ\mathsf{S}_{\alpha^{-\frac{1}{2}}}$, which is antisymmetric in the Frobenius norm (see lemma 6), and has better convergence behavior. The number of Taylor terms ($m^{*}$) and the scaling factor $(s^{*})$ are based on [2], as discussed in appendix A.
Algorithm 1 Parallel transport algorithm for Stiefel and flag manifolds
[…]
6:return $\begin{gathered}[Y|Q]\operatorname{expm}(t\begin{bmatrix}2\alpha A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})We^{t(1-2\alpha)A}+(\eta-YY^{\mathsf{T}}\eta-QQ^{\mathsf{T}}\eta)e^{t(1-\alpha)A}.\end{gathered}$
In table 2, we show the computational time along a time grid, for different $n$ and $\alpha\in\{\frac{1}{2},1\}$ for Stiefel manifolds, and for $\alpha=\frac{1}{2}$ for flag manifolds. The table generally shows sublinear growths in execution time in $n$ and $t$, with large $t$ behavior closer to linear. This is consistent with the $O(C_{1}nd^{2}+C_{2}td^{3})$ cost. table 3 shows execution time by $d$, also along the same grid. The cost is sub-cubical in $d$, and is closer to cubical as $t$ increases, again, consistent with a sum of $d^{2}$ and $d^{3}$ terms.
In fig. 1, we show the approximate $O(nd^{2})$ cost for small $t$ and larger range of $d$ and $n$ for Stiefel manifolds (the situation for flag manifolds is similar), with $d$ up to $500$ and $n$ up to $16000$. The graph shows a sub-quadratic cost (without the exponential action component) in $d$ (values at $d=100,200,400$ are $0.100,0.283$, and $0.882$). The cost in $n$ is also sublinear (values at $n=1000,2000,4000$, $8000$ and $16000$ are $0.0162,0.0203,0.0307,0.0458$, and $0.0802$, respectively). The fact that we do not get exactly quadratic and linear costs is likely due to fixed time cost incurred in the computer operation.
(a) Time versus $d$
(b) Time versus $n$
Figure 1. Execution time versus $n$ and $d$ for Stiefel manifolds. The dotted lines are times without the cost of the exponential action. The solid lines are with the exponential action. On the left, we fix $n=1000$ and let $d$ vary from 22 to $500$. The graph is approximately quadratic in $d$. On the right, we set $d=20$ and let $n$ vary from $70$ to $16000$. The graph has a linear trend in $n$. Note the gap due to the exponential action increases with $d$ but is relatively constant versus $n$, consistent with the cost $O(td^{3})$.
Using automatic differentiation, in [32] we can verify that these correspond to highly accurate computation, depending on the manifold, with the transport equation satisfied with absolute accuracy around $10^{-10}$ to $10^{-12}$.
Stiefel man. comp. time Flag man. comp. time.
Geodesic time grid $t$ Geodesic time grid $t$
$\alpha$ n 0.5 1.0 2.0 5.0 20.0 0.5 1.0 2.0 5.0 20.0
$\frac{1}{2}$ 100 0.05 0.10 0.15 0.25 0.69 0.06 0.08 0.15 0.22 0.57
200 0.08 0.12 0.17 0.32 0.80 0.09 0.13 0.17 0.30 0.75
1000 0.15 0.18 0.26 0.47 1.45 0.14 0.18 0.26 0.49 1.41
2000 0.18 0.24 0.34 0.59 1.78 0.14 0.25 0.32 0.54 1.93
1 100 0.09 0.16 0.22 0.50 1.15
200 0.12 0.17 0.25 0.45 1.44
1000 0.18 0.21 0.35 0.68 2.36
2000 0.21 0.28 0.41 0.85 2.88
Table 2. Computational time for parallel transport for Stiefel and flag manifolds by $n$ for $d=50$. The flags are (20,15,15,nd)(20,15,15,n-d). The canonical metric corresponds to $\alpha=\frac{1}{2}$, the Euclidean embedded metric corresponds to $\alpha=1$. The table shows sublinear growth in execution time as $n$ and $t$ increase. Execution is on a free colab session. All time units are in second(s).
Stiefel man. comp. time Flag man. comp. time.
Geodesic time grid $t$ Geodesic time grid $t$
$\alpha$ d 0.5 1.0 2.0 5.0 20.0 0.5 1.0 2.0 5.0 20.0
$\frac{1}{2}$ 5 0.01 0.02 0.04 0.08 0.25
10 0.03 0.06 0.10 0.15 0.43 0.03 0.05 0.08 0.15 0.43
20 0.05 0.08 0.13 0.21 0.53 0.05 0.07 0.15 0.21 0.54
100 0.42 0.60 0.99 2.55 9.18 0.45 0.57 1.13 2.35 8.60
1 5 0.01 0.02 0.04 0.10 0.37
10 0.04 0.07 0.13 0.23 0.61
20 0.07 0.10 0.15 0.25 0.73
100 0.60 0.92 1.68 3.92 15.80
Table 3. Computational time for parallel transport for Stiefel and flag manifolds by $d$ for $n=1000$. We scale the partial flag of $(d_{1},d_{2},d_{3})=(4,3,3)$ by $\frac{d}{4+3+3}$ to $d$ for flag manifolds, setting $d_{4}=n-d$. Computational time is subcubical in $d$, consistent with the cost $O(C_{1}nd^{2}+td^{3})$. For large $t$, the trend is approximately linear in $t$. All time units are in seconds.
10.1. Parallel transport as isometries
We also demonstrate numerically the isometric property of parallel transport as implemented by our method, by showing the inner product matrix of a set of vectors is preserved during transportation. We consider a geodesic starting at $\operatorname{I}_{2000,100}$, with twenty tangent vectors of integer lengths sampled from integers between $1$ and $60$, while the relative angles are randomly generated. The geodesic initial velocity vector $v$ has length $1$. We compute the inner product matrix of the random tangent vectors at $t=0$ and of the transported vectors along a time grid. We measure the maximum absolute difference between the inner product matrices at each grid point and the initial point and plot the $\log_{10}$ of the differences in fig. 2. The absolute errors are around $10^{-12}$ up to $t=15$.
(a) Stiefel manifolds
(b) Flag manifolds
Figure 2. Isometry of parallel transport of Stiefel and flag manifolds. We transport a set of $20$ vectors from an initial point on $\mathrm{St}_{2000,100}$, with $\alpha=1$ along a geodesic, and compute the maximum difference between the inner product matrices of these 2020 vectors at $t=0$ versus $t\in\{0.1,.3,.5,.7,1.2,1.5,1.7,2.1,3.,15.\}$. On the right-hand side, we consider $\operatorname{Flag}(50,20,30,1900)$ with the canonical metric $\alpha=\frac{1}{2}$. Up to $t=15$ seconds, the differences are in the order of $10^{-12}$.
11. Discussion
In this paper, we provide an efficient algorithm to compute parallel transport on Stiefel and flag manifolds, as well as transport formulas for the general linear group and the special orthogonal group. We also provide a framework for parallel transports on groups with a deformation metric based on a pair of Lie algebras with bi-invariant metrics, in particular, a pair of transposable Lie algebras. We believe the framework developed here will be useful for other manifolds in applied mathematics.
Acknowledgments
We wish to thank the two anonymous reviewers and the editor for the careful review and numerous suggestions that helped us improve the article significantly. In particular, we are grateful for one reviewer’s strong suggestion to use the reductive space approach, which pointed us in the direction of establishing the naturally reductive property. The new approach gives us a much clearer perspective of the transport formula, obtained by solving the transport equation directly in our first version. The same reviewer’s suggestion also leads to a better derivation of the Christoffel functions from the geodesic equations. We thank him/her for the references [35, 25] in relation to lemma 4 as well as the observation that the algorithm could be considered $O(nd^{2})$ for bounded time. Needless to say, any remaining mistake is our sole responsibility.
Du Nguyen would like to thank his family for their loving support in this project. He thanks Mr. Tom Szwagier for feedback on an earlier version of the article. The work was supported by a research grant (VIL40582) from VILLUM FONDEN, and the Novo Nordisk Foundation grants NNF24OC0093490 and NNF24OC0089608.
Appendix A Exponential action
The solution of a linear ODE with constant coefficient $\dot{X}=AX$ and initial condition $X(0)=X_{0}$ is given by $\operatorname{expm}(tA)X_{0}$, where typically $X$ is a vector and $A$ is a matrix, where $\operatorname{expm}$ is the matrix exponential. We often need to treat the case where $X_{0}$ is a matrix and $A$ is an operator on a matrix space. In that situation, which is the situation encountered in this article, computing the exponential by vectorization is usually inefficient. The better alternative is to compute the exponential action, denoted by $\operatorname{expv}(tA,X_{0})$. Instead of evaluating the full matrix exponential, we approximate the action $\operatorname{expm}(tA)X_{0}$ more efficiently, often [2] by evaluating a scaling and squaring of a Taylor expansion of the operator exponential operating on $X_{0}$.
Specifically, the popular library algorithm for exponential action in Python (function expm_multiplyexpm{\_}multiply in the scipy package [43] and similar functions for other languages, expmvexpmv in MATLAB and expvexpv in Julia) are based on the work of Al-Mohy and Higham [2], approximating $e^{A}B=(e^{1/sA})^{s}B\approx(T_{m}(s^{-1}A))^{s}B$, where $T_{m}$ denotes the $m$ terms Taylor series expansion of $e^{A}$. The cited paper provides an algorithm to find a pair of optimal $(m,s)$, based on estimates of $\|A^{p}\|^{1/p}_{1}$ for several values of $p$. The paper found the higher $p$ estimates are beneficial when $A$ is not normal, for example, when $A$ is upper triangular. These estimates are used to determine the Taylor order $m$.